OpenStudy (anonymous):

The solution of the equation 2/(5x+5)-3/(x^2-1)=4/(x-1)

hero (hero):

\[\frac{2}{5(x+1)} - \frac{3}{(x+1)(x-1)} = \frac{4}{x-1}\]

hero (hero):

Multiply both sides by x + 1 to get: \[\frac{2}{5} - \frac{3}{x-1} = \frac{4(x+1)}{x-1}\]

hero (hero):

Add 3/(x+1) to both sides to get: \[\frac{2}{5} = \frac{4(x+1)}{x-1} + \frac{3}{x-1}\]

hero (hero):

Combine fractions on the right side to get: \[\frac{2}{5} = \frac{4x + 4 + 3}{x-1}\] Simplify the numerator get: \[\frac{2}{5} = \frac{4x + 7}{x-1}\]

hero (hero):

Cross Multiply to get: 2(x-1) = 5(4x + 7)

hero (hero):

Expand both sides to get: 2x - 2 = 20x + 35 Place like terms on the same side to get -35 - 2 = 20x - 2x -37 = 18x

hero (hero):

Finish solving for x from there.

OpenStudy (anonymous):

are you going to divide 18 by 37

hero (hero):

No, you're going to divide both sides by 18 to isolate x

OpenStudy (anonymous):

-37/18